A machine that have the following cost is under consideration for a new manufacturing process. What is the equivalent annual worth? The MARR is 10% compounded semiannually. The first cost is $50,000, the semiannual operating cost is 10,000, the semiannual income is 20,000, the semiannual income gradient is 100, the salvage value is 5,000 and the life in years is 4 years a. EAW- -$4,112 b.EAW $2,112 G. EAW--$3,112 d. EAW -$3,112
Answer #1

Ans: EAW = $3112

Explanation:

Here, n = (12 / 6) * 4 = 8

          i = 10% / 2 = 5%

EAW = -50,000(A/P, 5%, 8) - 10,000 + [20,000 + 100(A/G, 5%, 8)] + 5,000(A/F, 5%, 8)

         = -50,000(0.1547) - 10,000 + [20,000 + 100(3.24)] + 5,000(0.1046)

         = -7,735 - 10,000 + 20,324 + 523

         = $3112